If you put $700 at the beginning of the 2nd month, in a savings account with a 10% nominal rate of interest compounded monthly, what will the investment be worth in 12 month (value at the end of the first year, round to the nearest dollar)?
Answer #1

Solution:-

The formula for calculating future value is as follows:

FV = PV*(1+r)n

where,

FV= Future value

PV= Initial investment

r= rate of return per compounding period

n= no. of compounding periods

In the given question, we have the following information provided:

FV= ?

PV= $700

r (return per month)= 10%/12 = 0.83%

n= 11 (i.e. 11 months since investment is made in the second month)

Hence,

FV= 700*(1+0.83%)11

FV= $767

Therefore, the investment will be worth $767 in 12 months time.